OpenStudy (nosgood1):

what's the exact distance from (-5, 1) to (3,0)

OpenStudy (sooobored):

use the distance formula \[d=\sqrt{(x_1-x_2)^2 +(y_1-y_2)^2}\]

OpenStudy (nosgood1):

(-5-3)= -5+3=8 + 1-0= 1 so 8+1=9 =3?

OpenStudy (will.h):

-5-3 is -8 not 8 besides don't forget about the exponent 2 (squareD)

OpenStudy (will.h):

So it would be sqrt65